OpenStudy (anonymous):

how do you evaluate f(x) at the given values of x f(x)=x^1/4 x=81,-3

OpenStudy (anonymous):

x=-3 is not in the domain of f

OpenStudy (anonymous):

f(81)=81^1/4=(3^4)^1/4=3

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = x^{\frac{1}{4}}\] \[\Large f(81) = 81^{\frac{1}{4}}\] \[\Large f(81) = \sqrt[4]{81}\] \[\Large f(81) = \sqrt[4]{3^4}\] \[\Large f(81) = 3\] As for the other value, you cannot take the fourth root of a negative number.